Try: circle area, U = I × R, radius
Equally likely outcomes, complements, sums and at least one event.
P(A)=m/n
P(Ā)=1−P(A)
P(A∪B)=P(A)+P(B)
P(X≥1)=1−(1−p)ⁿ
odds = p/(1−p)